### v1

# Can We Verify Step by Step for Incorrect Answer Detection?

Abstract
Chain-of-Thought (CoT) prompting has marked a significant advancement in enhancing the reasoning capabilities of large language models (LLMs). Previous studies have developed various extensions of CoT, which focus primarily on enhancing end-task performance. In addition, there has been research on assessing the quality of reasoning chains in CoT. This raises an intriguing question: Is it possible to predict the accuracy of LLM outputs by scrutinizing the reasoning chains they generate? To answer this research question, we introduce a benchmark, R2PE, designed specifically to explore the relationship between reasoning chains and performance in various reasoning tasks spanning five different domains. This benchmark aims to measure the falsehood of the final output of LLMs based on the reasoning steps. To make full use of information in multiple reasoning chains, we propose the process discernibility score (PDS) framework that beats the answer-checking baseline by a large margin. Concretely, this resulted in an average of $5.1\%$ increase in the F1 score across all 45 subsets within R2PE. We further demonstrate our PDS’s efficacy in advancing open-domain QA accuracy. Data and code are available at https://github.com/XinXU-USTC/R2PE.
Can We Verify Step by Step for Incorrect Answer Detection?
Xin Xu, Shizhe Diao, Can Yang, Yang Wang Hong Kong University of Science and Technology xxuca@connect.ust.hk {sdiaoaa, macyang, yangwang}@ust.hk
Figure 1: An example from HotpotQA that GPT-4 outputs highly consistent but wrong answer. In this particular example, our PDS can detect conflicting information about the actor (colored by red) and make a prediction that aligns with the label, while the ADS that uses the answer agreement does not.
footnotetext: †Corresponding author.
Figure 2: An overview of R2PE benchmark and the PDS framework. The construction of R2PE encompasesses 5 stages: data collection, response generation, extraction, manual cleaning, and labeling. Sourced from eight distinct datasets covering a variety of task types and derived from 6 different LLMs, R2PE comprises 45 subsets, featuring approximately 20.2K TRUE and 18.1K FALSE instances across 5 diverse domains. The objective is to establish a discernibility score that accurately indicates the veracity of the answers. PDS adopts implicit answer checking and explicit process supervision to detect all potential discrepancies among different reasoning pathways, which is superior to ADS that merely focuses on the answer consensus.
1 Introduction
Recent development in large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Bubeck et al., 2023) has showcased their remarkable aptitude for tackling diverse downstream tasks. Given several demonstrations with reasoning steps, LLMs exhibit a formidable capability to address reasoning tasks, commonly referred to as chain-of-thought (CoT) (Wei et al., 2022). There have been significant advancements in improving the reasoning abilities of LLMs in terms of end-task performance, such as intricate math word reasoning (Shum et al., 2023; Diao et al., 2023; Zheng et al., 2023; Zhang et al., 2023b), decision-making tasks (Yao et al., 2022), and tasks that involve extensive search and tactical planning (Yao et al., 2023; Besta et al., 2023; Ding et al., 2023). Another line of research lies in the analysis of the reasoning steps themselves with or without human-annotated ones (Clinciu et al., 2021; Welleck et al., 2022; Saparov and He, 2022; Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023).
Wei et al. (2022); Ye and Durrett (2022) have manually inspected whether reasoning steps aligns with the correctness of ultimate answers in both correct and incorrect answer groups through case studies. Moreover, Prasad et al. (2023); He et al. (2023) find that elevating the quality of the rationales could potentially enhance task performance. Since aggregate task performance is a summation of the accuracy of final answers across individual instances, these observations offer an initial qualitative scrutiny of the rationales and their impact on the accuracy of the final predictions. Despite these insights, there remains a shortage of quantitative proof to substantiate whether the evaluation of reasoning chains can reliably affirm the validity of the final outcomes. To bridge this gap, we introduce the R2PE (Relationship of Rationales to Performance Evaluation) benchmark, a test bed designed to quantitatively investigate this question.
The susceptibility of LLMs to generate incorrect information has been underscored by OpenAI-Blog (2022); Zhao et al. (2023a). Meanwhile, Ye and Durrett (2022) associate false predictions with nonfactual explanations. Furthermore, Wei et al. (2022) have demonstrated that, even in arithmetic reasoning tasks, incorrect rationales can occasionally yield correct outcomes, and a variety of errors may occur at intermediate reasoning steps. To perform a qualitative evaluation of rationales with the aim of validating final predictions, the R2PE benchmark is established to integrate a diverse spectrum of reasoning tasks, covering mathematical, commonsense, physical, and textual reasoning (including fact verification and open-domain question answering). These tasks extend across domains such as mathematics, common knowledge, physics, literature, and general world knowledge. Responses are collected from six distinct LLMs to promote a broad generality of the findings. The characteristics and the creation steps of the R2PE benchmark are illustrated in Figure 2.
Each evaluated question or claim begins with the generation of multiple rationales, leading to an answer aggregation based on majority voting (Wang et al., 2022) after extraction and manual cleaning. Once the final outcomes are labeled as true or false using an exact match, our goal is to derive a discernibility score (DS), which is intended as an indicator of the credibility of the final answer. To fully exploit the information contained in all reasoning chains, we present a process discernibility score (PDS) that substantially exceeds the answer discernibility score (ADS) baseline that counts the number of the same answers in terms of the F1 score. As shown in Figure 1, PDS can detect conflicting information about the actor and predict the incorrectness of the final answer. Figure 2 presents a succint overview of both our benchmark and the associated methodology.
In summary, our contributions are as follows:
  • We propose R2PE, the first benchmark that quantitatively assesses the relationship between reasoning chains and end-task performance across a spectrum of reasoning tasks, multiple domains, and an array of LLMs.
  • We introduce a process discernibility score (PDS) that aggregates the information in different chains of reasoning to validate the correctness of the final output.
  • Comprehensive experiments of PDS in R2PE reveal its superiority over the answer discernibility score (ADS), leading to consistent increases in the F1 score.
  • Further experiments show the effectiveness of our approach in improving open domain question answering performance when combined with verify-and-edit (Zhao et al., 2023a).
2 R2PE Benchmark
R2PE serves as a comprehensive platform for assessing the reliability of LLM-generated answers in CoT reasoning. R2PE is meticulously constructed with several critical characteristics in mind: (i) Assessability: every instance within R2PE can be verifiable as true or false based on a certain criterion. (ii) Task Diversity: The benchmark should encompass a wide range of reasoning datasets, featuring various answer formats, and spanning different task categories across multiple domains. (iii) Generalizability across LLMs: Responses should be elicited using different LLMs to ensure broad applicability. (iv) High Quality: To minimize instances where correct answers are inaccurately labeled as false due to extraction failures, it is crucial that answers are precisely extracted from responses.
2.1 Construction process
As shown in Figure 2, the creation of R2PE consists of the following five steps:
Data Collection. we utilize a total of eight datasets including benchmarks for mathematical reasoning such as GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021), common sense reasoning tasks like StrategyQA (Geva et al., 2021) and Play dialogue from BIG-bench collections (Srivastava et al., 2022), physical reasoning tasks (Physics from BIG-bench collections (Srivastava et al., 2022)); fact verification (FEVER (Thorne et al., 2018)) and open-domain question answering (HotpotQA (Yang et al., 2018), 2WikiMultihop (Ho et al., 2020)). Each dataset selected meets the criterion of assessability through verifiable answers. These tasks not only vary in answer formats, including numerical, yes/no, multiple choice, and free form, but also span an array of domains such as mathematics, commonsense, literature, physics, and general world knowledge (see Appendix A.1).
Response Generation. For each question (or claim in the FEVER dataset) $Q$, the LLM is prompted using CoT to produce $n$ responses $(R_{1},R_{2},...,R_{n})$, respectively. The full prompts are given in Appendix F. The responses are aggregated from six distinct LLMs to ensure the generalizability: text-davinci-003 (Ouyang et al., 2022), GPT-3.5-turbo (OpenAI-Blog, 2022), its instruct-trained variant GPT-3.5-turbo-instruct, Gemini Pro, Mixtral-8x7b (Jiang et al., 2024), and mistral-medium. A concise exposition on LLMs along with the particulars pertaining to each model, is deferred to Appendix A.2. Specifically, we set the temperature to 0.7, $n=5$ for all datasets and ’topK’ to be 40, ’topP’ to be 0.95 in the Gemini Pro API.
Extraction. Ratinale $r_{i}$ and the corresponding anwser $a_{i}$ need to be isolated from each response $R_{i}$. During the pilot extraction period, we use the prompt cue "The answer is" as a delimiter to segregate responses. Our observations indicated that almost all LLMs yielded some outputs that did not conform to the expected answer format. Given the varying responses styles across different LLMs, the answer trigger words are identified for different LLMs and datasets after observing the original responses to facilitate extraction (see Table 6). This strategy effectively segments most of the responses.
Manual Cleaning. To maintain a high-quality benchmark, manual inspection and cleaning are performed to handle unusual cases. For atypical responses that deviate from the recognized patterns, we either manually separate the response to derive the answer or assign a special marker when separation is not feasible. Additionally, special attention has been paid to non-sentical responses. A detailed description of the extraction and manual cleaning procedure is provided in Appendix A.3.
Labeling. Upon completion of extraction and manual cleaning, $n$ responses with reasoning chains and corresponding responses $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{n}=(r_{n},a_{n})\}$ are collected for each question $Q$. With these pairs of rationale answers extracted, we can aggregate the outputs to obtain the final answer by the majority votes: $a=argmax_{a_{i}}freq(a_{i})$, where $freq(a_{i})$ denotes the frequency in which $a_{i}$ appears. The final result $a$ of each question $Q$ is then cross-referenced with its ground truth; A match results in a "TRUE" label, while a mismatch is designated as "FALSE".
2.2 R2PE Overview
Our R2PE benchmark comprises a diverse collection of data derived from six LLMs on eight reasoning tasks. These tasks encompass a variety of types, answer formats, and domains, making the benchmark rich and comprehensive. Overall, it includes approximately 38.3K instances with around 20.2K labeled as TRUE and 18.1K as FALSE. The structure of R2PE allows for the organization of instances into 45 distinct subsets based on the original dataset and the LLM used to generate responses. Hereinafter, subsets are denoted as (<dataset name> from <LLM name>), with detailed statistics and concrete examples available in Appendix B.
Each instance $e$ within R2PE incorporates a series of structured data fields: question or claim $Q$, the associated dataset name, the queried LLM name, five responses alongside their reasoning chains and answers $R=\{R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})\}$, the final output $a$, the ground-truth answer of question $Q$, and the label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth. Detailed explications of the constituent data fields, accompanied by an illustrative example, are shown in Table 10. Additional examples are presented in Appendix B.
With R2PE, we aim to predict the correctness of final outcomes based on the intermediate reasoning steps. The response set $R$ serves as an input for predicting the label $L$ for every instance $e$. We introduce a numerical criterion, the discernibility score (DS), to encapsulate the quality of the rationales and their corresponding answers. A low DS might suggest a potential mismatch between the final output $a$ and the ground-truth answer. Hence, we will classify the example as false: $\hat{L}=F$, if its DS falls below a certain threshold $H$.
3 Process Supervision
Having R2PE at our disposal, our goal is to find a suitable DS that can be a good indicator for the veracity of final predictions. We suggest the process supervision to be taken into account recognizing its potential to overcome the limitations of DS based solely on answer agreement, which tends to overlook a great number of examples with the label $L=F$. To demonstrate its usefulness, we have conducted comprehensive experiments (results are presented in Section 4).
3.1 The Fallacy of Answer Agreement
Based on the insights of Wang et al. (2022); Zhao et al. (2023a), the degree of consensus among answers, is posited as a diagnostic tool to gauge the instances in which LLMs may make wrong predictions. We refer to the answer aggreement number as Answer discernibility Score (ADS) hereinafter. The ADS can thus be adopted as a natural baseline in our R2PE benchmark. A high ADS often suggests a high likelihood that the proposed answer is correct. Given this premise, designating a threshold of $H=n/2$ is a reasonable strategy. If the ADS is below that midpoint, $\hat{L}=F$ is predicted.
However, LLMs may generate substantially concordant but incorrect answers in reasoning tasks. Figure 1 showcases one such example, where the LLM is queried about the nationality of an actor from "Hostage for a Day". The LLM gives the answer "American" four times and "Canadian" once. Here, the ADS stands at 4, which erroneously points to a seemingly accurate but ultimately false consensus. This example serves to warn against overreliance on ADS as the sole indicator of the veracity of the final answer in reasoning tasks.
Figure 3: Qalitative analysis of results on the subset (HotpotQA from GPT-4): ADS (left), PDS (middle), and PDS-ADS (right). ADS has high precision: large answer agreement does not always guarantee accurate predictions. PDS yields desirable outcomes, while PDS-ADS has poor precision.
3.2 Process Supervision
The ADS appears to be an inadequate indicator of the ground truth label $L$, a limitation that may be ascribed to its myopic focus on the answer aggreement, while disregarding potential inconsistencies within the reasoning chain, as indicated by the areas marked red in Figure 1. Each response $R_{i}$ constitutes a rationale $r_{i}$ paried with an answer $a_{i}$. The ADS, however, limits its analysis to the answers $(a_{1},...,a_{n})$, which takes up only a small proportion of the information contained in $R$. Considering the demonstrated significance of intermediate reasoning steps (Yoran et al., 2023), a DS that makes use of rationales would be preferable.
Inspired by the hypothesis that LLMs tend to generate inaccurate or irrelevant information when they are uncertain about their predictions (Manakul et al., 2023), we introduce a process discernibility score (PDS). The PDS is designed to assess potential discrepancies among all responses, going beyond explicit verification of answers alone.
As illustrated in Figure 2, our PDS is defined as
$PDS=(\overline{ADS}+PSS)/2,$ (1)
where $\overline{ADS}=2(ADS-n/2)/n$ is a normalized ADS, which adjusts the original ADS to a balanced scale around zero. The Process Supervision Score $PSS=mean_{i\neq j}PPSS(R_{i},R_{j})$ is the average of all Pairwise PSS (PPSS). $PPSS(A,B)$ can be any score that quantifies the extent to which the content of B is included in and contradicted by A.
Subsequently, in a manner analogous to the ADS prediction procedure, we classify instances where the PDS falls below the threshold $H$ as $\hat{L}=F$ (indicating a likely falsehood), whereas all remaining instances are designated as $\hat{L}=T$.
Compared to ADS, our PDS is capable of overseeing reasoning processes as well as implicit answer checking. We would certainly expect a performance improvement over the ADS (see Section 4). As depicted in Figure 1, the PDS can detect contradictive information about the name of the actor playing a Russian hostage taker in "Hostage for a Day” among multiple reasoning chains (highlighted by red) and give a correct label $\hat{L}=F$, while the ADS only assesses the agreements among the answers and leads to a wrong prediction $\hat{L}=T$.
Dataset Method GPT3 GPT-instruct GPT-3.5 Gemini Mixtral mistral avg
GSM8K ADS 66.59 53.63 42.83 52.05 58.63 50.14 -
PDS 69.30 56.65 52.50 55.04 62.50 53.26 +4.23
MATH ADS 81.33 76.72 70.40 77.92 81.19 77.56 -
PDS 86.55 80.38 74.93 78.95 83.85 81.33 +3.48
StrategyQA ADS 36.79 59.06 54.28 52.54 60.39 62.59 -
PDS 52.21 59.14 57.51 56.76 63.14 63.83 +4.49
Play ADS 50.22 56.64 54.28 53.07 71.12 56.32 -
PDS 55.59 59.00 56.16 54.09 72.02 58.78 +2.30
Physics ADS 48.70 52.46 56.52 65.86 61.54 63.33 -
PDS $\textbf{52.31}^{b}$ 55.56 58.65 88.44 66.90 66.01 +6.58
FEVER ADS 50.92 55.97 49.14 63.56 64.58 - -
PDS 59.64 60.68 53.01 $\textbf{63.64}^{b}$ 65.72 - +3.70
HotpotQA ADS 72.78 71.52 $63.95^{a}$ 81.46 74.24 - -
PDS 85.71 78.84 $\textbf{70.25}^{a}$ 85.65 78.66 - +7.00
2WikiMultihop ADS 69.26 69.75 $42.67^{a}$ 62.83 69.36 - -
PDS 78.65 76.51 $\textbf{57.14}^{a}$ 71.76 70.81 - +9.80
avg - +7.92 +3.88 +5.76 +5.63 +2.82 +2.65 +5.10
Table 1: PDS consistently outperform ADS across all subsets in our R2PE benchmark. All metrics are F1 score (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively. a. These four figures are from the (HotpotQA from GPT-4) and (2WikiMultihop from GPT-4) respectively (see Section 2). b. These two are reported using a different threshold, and the details are discussed in Appendix C.
3.3 Experiment Setup
Experiments are carried out to show the effectiveness of PDS across all subsets of the R2PE benchmark. We will employ the F1 score as our evaluation metric (see Section 2), with precision and recall also included as supplementary measures. We establish the ADS as our baseline. For subsets comprising binary (yes/no) choices or multiple-choice queries, there are inherently limited answer possibilities, which often results in greater answer agreement among responses. To accommodate this characteristic, the threshold $H$ for ADS is set to 4.5 for subsets in R2PE associated with StrategyQA, Play, Physics, and FEVER. On the other hand, in subsets originated from GSM8K, MATH, HotpotQA, and 2WikiMultihop, the midpoint $H=2.5$ is chosen.
For the $PPSS(\cdot,\cdot)$ score, we adopt the SAUMMAC zero-shot model (Laban et al., 2022) that trained on MNLI (Williams et al., 2018). It is a consistency detection method for text summarization that leverages the out-of-the-blox natural language inference model to compute pairwise entailment score. As suggested by Laban et al. (2022), the probability of sentence-level entailment and contradiction is used. To detect all potential inconsistency in any sentence, we use $min$ operation to aggregate sentence-level entailment and contradiction:
$PPSS(A,B)$ $=\min_{j}\max_{i}\bigl{(}ent(a_{i},b_{j})$ (2)
$\quad-con(a_{i},b_{j})\bigr{)},$
where $a_{i},b_{j}$ are sentences of $A,B$ correspondingly, $ent$ represents the probability of entailment, $con$ is the probability of contradiction, the $min$ is taken over all sentences of $B$, and $max$ is taken over all sentences of $A$. The range of this PPSS spans in $(-1,1)$. A negative value implies the potential presence of confilcting information. Consequently, it is reasonable to set the threshold $H$ for the PDS to be 0 for numeric and free form questions and $0.4$ otherwise. Comprehensive ablation studies pertaining to alternative PPSS choices are discussed in Section 4.2. The F1 score is employed to assess the performance in the various subsets of R2PE.
4 Results and Analysis
We present our findings in Table 1, which shows that PDS outperforms the baseline in all subgroups of the R2PE benchmark. To investigate the effects of the components of PDS (see Equations (1) and (2)), ablation studies are carried out in subsets related to HotpotQA and 2WikiMultihop. Finally, we demonstrate the effectiveness of our approach in enhancing the accuracy of open-domain question answering by integrating it with the verify-and-edit (VE) framework (Zhao et al., 2023a).
Method GPT3 GPT-instruct GPT-4 Gemini Mixtral
PDS 85.71 78.84 70.25 85.65 78.66
PDS - ADS 82.00 78.16 60.57 82.35 74.24
PDS w/o ans 78.44 78.84 63.41 84.79 74.19
PDS-avg 77.49 78.81 62.72 84.86 74.29
PDS-Halocheck 77.49 78.11 62.72 84.99 75.32
PDS-selfcheckNLI 45.77 51.01 33.62 66.87 60.25
PDS 78.65 76.51 57.14 71.76 70.81
PDS - ADS 76.97 73.35 55.00 70.14 69.36
PDS w/o ans 76.12 75.56 50.19 67.38 68.44
PDS-avg 76.95 75.34 51.06 67.23 68.87
PDS-HaloCheck 76.66 75.08 52.30 68.39 68.78
PDS-SelfCheckNLI 52.61 47.65 25.64 42.89 55.70
Table 2: Ablation study of PDS. All metrics are F1 score (in %). Top: subsets related to HotpotQA on R2PE. Bottom: subsets related to 2WikiMultihop on R2PE.
4.1 PDS substantially outperforms ADS
As shown in Table 1, across all 45 subsets in our R2PE, PDS consistently improves ADS performance, yielding an average improvement of $5.10\%$ in terms of F1. Although ADS demonstrates greater precision, it suffers from significantly lower recall, indicating its propensity to overlook numerous examples with $L=F$. In contrast, PDS adeptly balances precision and recall, culminating in superior overall performance relative to ADS (Table 14 and 15). This trend can be observed from Figure 3.
PDS demonstrates a stronger capability to verify the accuracy of the predictions for free-form questions than ADS, with notable average improvements of $7.00\%$ in HotpotQA and $9.80\%$ in 2WikiMultihop. This superior performance can be attributed to the process-oriented verification approach employed by PDS. By addressing the challenges associated with indirectly assessing answer consistency in free-form questions, PDS proves to be more effective in ensuring accuracy. It is important to note that the average improvement metrics provided in Table 1 serve primarily as an indicator of the extent to which PDS enhances performance over ADS. However, a direct comparison of F1 scores between different subsets within R2PE is not recommended. This is because the efficacy of the F1 score as a comparative metric is limited by variations in the ratio of TRUE and FALSE instances across different subsets.
To meticulously evaluate the efficacy of the PDS, we delve into the (HotpotQA from GPT-4) subset, partitioning it into two sets: one where answers fail to reach a consensus greater than half, and its counterpart. We then label the examples with an ADS less than the threshold $H$ as $\hat{L}=F$, and all the others as $\hat{L}=T$. As shown in Figure 3 (left), while the ADS proficiently segregates the examples with $L=T$, its discernment for the "False" category is less acute. On the contrary, the PDS (Figure 3 middle) strikes a better balance between precision and recall. Eliminating the implicit answer verification component in Equation (1) results in diminished precision (Figure 3 right, discussed in Section 4.2).
4.2 Ablation Study
To explore alternatives to PDS and understand the contribution of its components to the overall performance, we have carried out an extensive ablation study involving the following variants:
[…]
PDS w/o ans. The original PSS considers all $PPSS(R_{i},R_{j})$, where each response $R_{i}$ equals a reasoning path $r_{i}$ and an answer $a_{i}$. As the information from the answers $a_{i}$’ is already used to obtain the final answer $a$, we can discard them and only compute the entailment score among rationale pairs $PPSS(r_{i},r_{j})$, that is, $PDS_{w/o\>ans}=(\overline{ADS}+\sum_{i\neq j}PPSS(r_{i},r_{j}))/2$.
PDS-avg. We can change the aggregation operation $min$ of $PPSS(\cdot,\cdot)$ (Equation (2)) by taking the average. Then $PPSS(A,B)=\text{avg}_{j}\text{max}_{i}(ent(a_{i},b_{j})-con(a_{i},b_{j}))$.
PDS-HaloCheck;PDS-SelfCheckNLI. HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) can be alternatives to PSS, denoted by PDS-halocheck and PDS-selfcheckNLI, respectively. Discussion of the difference between these two against our PSS implementation is provided in Appendix C.4.
As shown in Table 2, our PDS implementation gains the best results.
4.3 PDS for Improving Downstream Performance
VE (Zhao et al., 2023a) post-edits reasoning process using external knowledge (Appendix D). VE first checks the answer agreement and then edits the reasoning chains based on a retriever if $ADS<H$. That is to say, VE selects potential incorrect samples by ADS. Our PDS can be incorporated into the VE framework, dubbed "VE + PDS".
Experiments are carried out on (HotpotQA from GPT-4) and (2WikiMultihop from GPT-4), utilizing ground-truth supporting contexts along with distractor paragraphs as the retrieved documents to negate the influence of the retrival system. Our baseline is VE, and the standard prompting method (Brown et al., 2020) and CoT-SC (Wang et al., 2022) are also included as a reference.
Figure 4 illustrates that integrating our method with the VE framework leads to enhanced accuracy for both datasets. Specifically, VE+PDS leads to a notable $2.9\%$ absolute increase in accuracy for HotpotQA, and $2.8\%$ for 2WikiMultihop. Given that the retrieval system in our experiments is assumed to operate with near-perfection, the observed improvements in performance are largely attributed to the more effective selection of samples, achievable through the substitution of ADS with PDS. This success is likely a consequence of the PDS capability to supervise processes across multiple reasoning chains, leading to better selection of all potential incorrect predictions.
Figure 4: PDS can be integrated with verify-and-edit to further improve accuracy on open-domain QA.
5 Related Work
5.1 Extensions of Chain-of-Thought
Chain-of-Thought (CoT) (Wei et al., 2022) encourages LLMs to detail their reasoning in natural language steps within few-shot examples. Wang et al. (2022) advance CoT by substituting greedy decoding with sampling techniques. Furthermore, various strategies (Chen et al., 2022; Gao et al., 2023; Zhang et al., 2023a) resort to prompting LLMs to express intermediate steps as Python codes. Given the sensitivity of few-shot learning to examplars, there has been an influx of research aiming to source high-quality demonstrations, from automatic annotations (Kojima et al., 2022; Zhang et al., 2022; Fu et al., 2022; Wan et al., 2023; Xu et al., 2023; Wang et al., 2023; Zhao et al., 2023b; Crispino et al., 2023) to hybrid approaches that blend manual design and automatic construction (Shao et al., 2023; Shum et al., 2023; Diao et al., 2023; Zou et al., 2023). For example, zero-shot-CoT (Kojima et al., 2022) proposes a bifurcated prompting approach; Fu et al. (2022) recommend selecting questions that require more reasoning steps while Diao et al. (2023) suggest selecting the most uncertain questions as demonstrations.
To enhance LLM reasoning across various tasks, novel prompt adaptations (Yao et al., 2022, 2023; Besta et al., 2023; Ding et al., 2023; Zheng et al., 2023; Zhang et al., 2023b) have been developed. ReAct (Yao et al., 2022) synergize reasoning with decision-making actions, while structural variants (Yao et al., 2023; Besta et al., 2023; Ding et al., 2023) tackle tasks requiring elaborate exploration and strategic planning. For complex mathematical problems and knowledge-intensive tasks, PHP (Zheng et al., 2023), cumulative reasoning (Zhang et al., 2023b), and verify-and-edit (VE) (Zhao et al., 2023a) have been crafted. Li et al. (2023) further extend VE framework by dynamically incorporating diverse knowledge sources.
Most prior research focuses predominantly on improving final task performance in terms of metrics, such as accuracy or exact match rates. Although a subset of them offers error analysis or failure cases, none of them extensively explores the relationship between the reasoning process and the resultant predictions. To the best of our knowledge, our work is the first to meticulously investigate whether the veracity of final predictions can be predicted by assessing the reasoning chains.
5.2 Reasoning Chain Quality Evaluation
Analyzing rationales is crucial for understanding the capabilities, limitations, and identifying refinement opportunities (Golovneva et al., 2022; Prasad et al., 2023; He et al., 2023). Methods to evaluate these chains fall into two categories: reference-based, which compares generated rationales to human-annotated "gold standards" (Clinciu et al., 2021; Welleck et al., 2022; Saparov and He, 2022), and reference-free approaches, like ROSCOE (Golovneva et al., 2022), which assess rationales using specific metrics without human-created references. Further advancements include fine-tuning and prompting techniques by Prasad et al. (2023) and He et al. (2023), respectively. Although Prasad et al. (2023); He et al. (2023) show that high-quality reasoning pathways can improve task performance, the direct link between reasoning chain quality and prediction correctness remains unexplored. He et al. (2023) qualitatively analyze reasoning chains based on prediction accuracy, but our work presents a comprehensive quantitative analysis to investigate this connection.
6 Conclusion
To shed more light on the connection between the quality of the reasoning steps and the final outcome, we present R2PE, the first benchmark that quantitatively analyzes whether we can validate the answer veracity by evaluating the reasoning chains on a variety of reasoning tasks across five different domains. Six different LLMs are used in the creation of R2PE to guarantee generalizability of our findings. We develop the PDS framework by ingegrating multiple rationales’ information to predict the falsehood of the final predictions, which significantly boosts the performance of answer checking on our R2PE benchmark. Furthermore, our approach can be easily combined with verify-and-edit framework to improve end-task performance.
Limitations
There are several limitations of our work. First, R2PE focuses exclusively on tasks with answers derived from a predefined set of options. Consequently, it does not cover open-ended generative tasks such as text summarization. Second, the creation of R2PE is based on the original CoT-SC implementation (Wang et al., 2022), which now has a plethora of extensions. Third, our PDS implementation only evaluates the relation among different reasoning chains, neglecting the internal logical consistency of individual chains. To address these limitations, future endeavors will involve the collection of additional examples that cover a broader spectrum of task types into R2PE, including those that require long-text generation. Moreover, we intend to use various extensions of self-consistency methods to enrich the R2PE benchmark. In addition, a more holistic assessment of the quality of reasoning chains is left for future work, and it is possible to integrate our method with other more advanced rationale editing techniques.
Ethics Statement
Our R2PE is designed as a benchmark for assessing the falsehood of final outputs produced by LLMs across various domains and different LLMs. We will make R2PE available under MIT license.
Our PDS offers a novel approach to evaluating the answer veracity of LLMs by examining the consistency among multiple reasoning chains. The PDS is computationally efficient and can be integrated with additional error-correction methodologies to enhance task-specific performance. On the one hand, the ability to flag inaccurate LLM predictions will prove invaluable to users, offering insights into the reliability of LLM-generated content. Additionally, once an answer is identified as incorrect, additional methods can be used for correction or further scrutiny, thus mitigating the risk of hallucination or misinformation.
Our work represents a foundational step towards cultivating more advanced methods that both high precision and high recall. Our effort not only enhances and reliability of LLMs, but also significantly contributes to the broader discourse on ethical AI development and deployment.
Appendix A R2PE Construction
This appendix incorporates details of the construction process of our benchmark R2PE, including dataset processing (Appendix A.1), queried LLMs (Appendix A.2), and details of manual cleaning (Appendix A.3).
A.1 Dataset Processing
We use eight datasets to create R2PE:
  • GSM8K (Cobbe et al., 2021) contains a collection of grade school math word problems. Our research uses the entire test set. MIT license: GSM8K license.
  • MATH (Hendrycks et al., 2021) consists of competition-level mathematics problems. For our study, we conducted a random sampling of 1,000 problems from its test set, which contains a total of 5,000 examples. MIT license: MATH license.
  • StrategyQA (Geva et al., 2021) is a question answering benchmark that focuses on open-domain questions where the required reasoning steps are implicit in the question and should be inferred using a strategy. We use the open-domain setting (question-only set) from BIG-bench Collaboration (Srivastava et al., 2022), and randomly select 1000 examples from its test set. MIT license: StrategyQA license.
  • Play originates from BIG-bench (Srivastava et al., 2022), designed to assess the model’s ability to deduce whether two closely sequenced "lines" from a dialogue in a Shakespearean play were spoken by the same character or by different characters. This particular task challenges the model’s understanding of context, character consistency, and dialogue progression within the intricate settings of Shakespearean works. We used the entire test set. Link: Play link.
  • Physics (from BIG-bench (Srivastava et al., 2022)), is crafted to evaluate a model’s understanding of physics by requiring it to identify the appropriate formula for solving physics word problems. This dataset comprises 229 multiple-choice questions. For our study, we used the first two questions as few-shot demonstrations, and the remaining questions serve as the test set. Link: Physics link.
  • The fact verification dataset, FEVER (Thorne et al., 2018), categorizes claims as "SUPPORTS," "REFUTES," or "NOT ENOUGH INFO" by analyzing evidence paragraphs obtained from Wikipedia. The data annotations of FEVER utilize information from Wikipedia, which is obtained in accordance with the Wikipedia Copyright Policy. We use the FEVER subset by Zhao et al. (2023a), where 1,000 of 3,000 samples are randomly selected to balance the number of right and wrong predictions in CoT.
  • Adversarial HotpotQA (Yang et al., 2018) is a multi-hop question answering dataset. We adopt the subset proposed by Ye and Durrett (2022), which presents a challenging set of data in which the correct and incorrect predictions are evenly balanced through the use of their model.
  • 2WikiMultihop (Ho et al., 2020) is a dataset designed for multi-hop question-answering tasks. It leverages the structured format found in Wikidata and utilizes logical rules to provide answers. It is released under the Apache License 2.0, allowing users to modify and use it for research purposes. We use the subset provided by Zhao et al. (2023a), which is a subsample with size 1000 out of the dev set of 12,576 samples.
As listed in Table 3, these tasks not only vary in answer formats, including numerical, yes/no, multiple choice, and free form, but also span an array of domains such as mathematics, commonsense, literature, physics, and general world knowledge.
We exclude open-ended generation tasks, such as text summarization, from our R2PE benchmark. These tasks can use proxy metrics (Lin, 2004; Ganesan, 2018; Yuan et al., 2021) to measure performance; however, it is difficult to label responses as true or false directly. A potential solution is to calculate the Pearson correlation between metrics and scores that capture the quality of reasoning chains, and we leave this to future work.
Dataset Task Type Answer Format Domain
GSM8K Mathematical Reasoning Numeric Mathematics
MATH Mathematical Reasoning Numeric Mathematics
StrategyQA Common Sense Reasoning Yes/No Commonsense
play Common Sense Reasoning Yes/No Literature
physics Physical Reasoning Multiple Choice Physics
FEVER Fact Verification Yes/No World Knowledge
HotpotQA Open-Domain QA Free Form World Knowledge
2WikiMultihop Open-Domain QA Free Form World Knowledge
Table 3: Summary of reasoning tasks, their answer formats, and domains in our R2PE benchmark.
A.2 Queried LLMs
We engaged six different LLMs to collect responses across various tasks to ensure generalizability.
  • GPT-3 (Brown et al., 2020), where we specifically use the text-davinci-003 engine. This engine became deprecated on 1 January 2024.
  • GPT-3.5-turbo-instruct, which demonstrates similar capabilities to models from the GPT-3 era. Compatible with legacy Completions endpoint and not Chat Completions, serving as an alternative to the text-davinci-003 engine.
  • GPT-3.5-turbo represents an advancement within the GPT-3.5 family, exhibiting the ability to understand and generate both natural language and code. We used the gpt-3.5-turbo-1106 engine.
  • Gemini-Pro, as released by Google, is positioned as the premier model in terms of scalability across an extensive array of tasks. The API is available at https://cloud.google.com/vertex-ai/docs/generative-ai/model-reference/gemini.
  • Mixtral-8x7B, a Sparse Mixture of Experts (SMoE) language model, was released by MISTRALAI. The model, characterized by its unique architecture and substantial scale, has claimed to outperform Llama 2 70B on most of the benchmarks it was tested against. It is available at https://huggingface.co/mistralai/Mixtral-8x7B-v0.1.
  • The mistral-medium model is an internal prototype model provided by MISTRALAI. The API is available at https://docs.mistral.ai/platform/endpoints/.
When examining performance in open-domain question-answering tasks, we replaced GPT-3.5-turbo with GPT-4 (OpenAI, 2023). The reason for this replacement was the observed reluctance of GPT-3.5-turbo to engage with questions from these datasets, often defaulting to responses indicating a lack of sufficient information. Attempts to solicit responses from the GPT-3.5-turbo model frequently resulted in non-committal replies such as ’I do not have enough information,’ also documented by Zheng et al. (2023) and illustrated through examples in Appendix E. We hypothesize that this behavior may be attributable to the removal of contexts, a characteristic of the open-domain nature of these tasks used in our study.
GPT-4 is recognized as one of the most advanced systems available, with our research utilizing the gpt-4-0314 engine during the data collection phase. All OpenAI APIs are available at https://platform.openai.com/docs/models/overview.
A.3 Extraction & Manual Cleaning
Following the query process, we set out to extract answers and rationales from the raw responses. Despite utilizing the answer trigger phrase ’The answer is’ within the prompt to facilitate the extraction of answers, numerous responses deviate from this expected format. Considering the variability in answer formats among datasets and LLMs, we focused primarily on segmenting responses that exhibited a certain pattern. Subsequently, we conducted meticulous manual inspections to address failure cases that do not follow the recognized pattern, thereby ensuring the quality of our R2PE benchmark. Here, we detail this procedure.
In the (GSM8K from text-davinci-003) subset, it is observed that for responses which do not follow the instructions, extracting the last numerical figure from the responses produces the answer accurately in most cases. It is important to distinguish between a full stop and a decimal point when the number is situated at the end of a sentence, as illustrated in Table 4. Nevertheless, there are scenarios where the final number does not represent the correct answer to the question, necessitating manual review and refinement. Of the (GSM8K from text-davinci-003) subset, a total of 8 cases required such manual intervention, some of which are detailed in Table 5. For these cases, we manually annotate the answers and revise the final output and the lable $L$ if needed.
Q: Jame will turn 27 in 5 years. In 8 years his cousin will be 5 years younger than twice his age. How many years separate the age of the two now?
A: Jame is 27 now, so in 8 years he will be 35. His cousin will be 5 years younger than twice his age, which would be 70.
Q: You can buy a movie super ticket for \$20 that includes rights to watch any movie and you get one popcorn with a soda for only \$1 extra. Or you can buy a movie ticket for \$12 and buy one popcorn and a soda at regular price. What’s the regular price of popcorn if the soda costs \$3 and you didn’t buy the super ticket and ended up saving \$2?
A: If the regular price of popcorn is x, then the total cost of a movie ticket and one popcorn and a soda is 12+x+3 = 15+x. If you bought the super ticket for \$20 and got one popcorn and a soda for \$1 extra, the total cost is 21. So 20 = 15+x. Solving this equation gives x = 5. The regular price of the popcorn is \$5.
Table 4: Examples of last number-based answer extraction from responses on GSM8K dataset.
Similar to the approach employed for the (GSM8K from text-davinci-003) subset, the strategy of extracting the last number is also applicable to the (GSM8K from GPT-3.5-turbo-instruct) subsets as well. This method results in 17 cases that do not conform to the regular pattern and thus require atypical handling. Some of these instances are documented for reference in Table 5.
For GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium, the responses are initially segmented with the phrase ’The answer is’. After that, the specific answer trigger words for each LLM are determined and used to further split the responses. These identified patterns are summarized in Table 6. This extraction step effectively processes most of the responses. Then the last number extraction is applied to the remaining responses. Subsequently, a meticulous manual review is conducted to address any peculiar cases. Specifically, on (GSM8K from Gemini Pro), there are only 14 atypical cases. These responses are manually corrected, with representative examples detailed in Table 5. Nevertheless, for the (GSM8K from GPT-3.5-turbo), (GSM8K from Mixtral-8x7b), and (GSM8K from mistral-medium) subsets,a greater number of unusual instances are encountered. This may be attributed to the fact that these LLMs do not follow clear instructions as well as others. To streamline the process, we employ heuristics for filtering certain atypical questions or claims where over 3 out of 5 responses fail to be segmented correctly by the phrase ’The answer is’ or through the identified answer trigger words (see Table 6).
(GSM8K from text-davinci-003)
Q: Sandra, the florist around the corner, is very unhappy with Fred’s incomplete order delivery. Sandra had ordered four times as many red roses as white carnations. She also ordered 200 pink calla lilies, which were five times the number of white carnations. Sandra has threatened to switch suppliers if the missing red roses are not delivered by 5 pm. To keep Sandra’s business, how many red roses must Fred deliver by 5 pm?
A: First, Sandra ordered 4 times as many red roses as white carnations. Second, she also ordered 200 pink calla lilies, which were 5 times the number of white carnations. Third, 4 * (200/5) = 160. Therefore, Fred must deliver 160 red roses by 5 pm to keep Sandra’s business.
Q: A simple folding newspaper or tabloid can be made by folding a piece of paper vertically and unfolding. Then, say, page 1 is printed on the left back, page 2 is printed on the left front, and then, perhaps page 32 is printed on the right back, and page 31 is printed on the right front. How many pieces of paper would be used in a 32-page tabloid?
A: For a 32-page tabloid, only one piece of paper is used.
(GSM8K from GPT-3.5-turbo-instruct)
Q: Sandra, the florist around the corner, is very unhappy with Fred’s incomplete order delivery. Sandra had ordered four times as many red roses as white carnations. She also ordered 200 pink calla lilies, which were five times the number of white carnations. Sandra has threatened to switch suppliers if the missing red roses are not delivered by 5 pm. To keep Sandra’s business, how many red roses must Fred deliver by 5 pm?
A: First, Sandra ordered 4 times as many red roses as white carnations. Second, she ordered 200 pink calla lilies, which were 5 times the number of white carnations. Third, to find the number of white carnations, divide 200 by 5, which equals 40. Fourth, to find the number of red roses, multiply 40 by 4, which equals 160. Fifth, Fred needs to deliver 160 red roses by 5 pm to keep Sandra’s business.
Q: Bahati, Azibo, and Dinar each contributed to their team’s 45 points. Bahati scored the most points and it was 20 more than Azibo scored and 10 more points than Dinar scored. How many points did Azibo score?
A: First, Bahati, Azibo, and Dinar each contributed 45 points. Second, Bahati scored 20 more points than Azibo and 10 more points than Dinar. Third, if we take away 10 points from Bahati’s score, we are left with 45 - 10 = 35 points. Fourth, if Azibo scored 35 points, then Bahati must have scored 35 + 20 = 55 points. Fifth, if Bahati scored 55 points and Azibo scored 35 points, then Dinar must have scored 45 - (55 + 35) = -45 points. Sixth, since Dinar cannot score negative points, we know that there must be a mistake in the problem. The answer cannot be determined.
(GSM8K from Gemini Pro)
Q: Benny saw a 10-foot shark with 2 6-inch remoras attached to it. What percentage of the shark’s body length is the combined length of the remoras?
A: First, the combined length of the remoras is 2 * 6 = 12 inches. Second, 1 foot is 12 inches. Third, 12 inches is 1 foot. Fourth, 1 foot is 1/10 of the shark’s body length.
Q: How many girls are in the school if 40% of a school population is made up of 240 boys?
A: First, 40% of the school population is made up of 240 boys. Second, to find the total population, divide 240 by 40% (or 0.4). Third, if 60% of the population is girls, multiply the total population by 60% (or 0.6).
Table 5: Examples of atypical cases that require manual cleaning on GSM8K dataset.
LLMs
GSM8K
StrategyQA
Play
Physics
text-davinci-003
Therefore, the answer is …
GPT-3.5-turbo-instruct
Therefore, …; So, the answer is …
Therefore, the answer is …
Therefore, the answer is …; Therefore, the correct answer is …
GPT-3.5-turbo
So, …; Therefore, …
Therefore, …; So, the answer is …
Therefore, the answer is …
Therefore, the answer is …
Gemini Pro
So, …; Therefore, …; Answer: …; The answer: …
Therefore, …
Therefore, …; …, the answer is …
Therefore, …
Claude-2
Therefore, …
Therefore, …; So in summary, …; So in conclusion, …; So the answer is …
Therefore, …; …, the answer is …
Therefore, the answer is …; So the answer is …; So the formula required is …
Mixtral-8x7b
Therefore, …; So, …
So, the answer is …; Therefore, the answer is …
Therefore, …; In conclusion, …; So, the answer is …; Thus, the answer is …
Answer: …; Therefore, …; So the answer is …; So, the answer is …; The final answer is …
mistral-medium
Therefore, …; So, …
Answer: …; Therefore, …
Answer: …; Therefore, …
Therefore, …; So, the answer is …; Answer: …
Table 6: Indentified answer trigger words on various subsets of the R2PE benchmark. The MATH dataset is not included because we utilize \$\textbackslash boxed\{\}$ to parse the answers without the need of these words.
LLMs GSM8K MATH StrategyQA Play
GPT3 8 1 0 0
GPT-3.5 17 1 0 0
GPT-3.5 69 0 0 0
Gemini 14 2 18 16
Mixtral 41 2 2 0
mistral 6 6 0 2
Table 7: The number of atypical cases across subsets of the R2PE benchmark. Noting that the numbers in the MATH, StrategyQA, and Play dataset indicate the counts of non-sentential responses.
In the MATH dataset, the answers are clearly demarcated with \$\textbackslash boxed\{\}$, which facilitates straightforward automated extraction. Specifically, we use the Python codes provided by (Zheng et al., 2023) to parse the answers. Two questions within (MATH from text-davici-003) and one quesion in (MATH from Mixtral-8x7b) are discarded due to parsing errors induced by the aforementioned code. In addition, a thorough manual inspection is conducted to address and correct the formatting in some unconventional cases. This was particularly necessary for certain questions where the responses deviated from standard sentential structures. The number of such cases is given in Table 7. Some of these examples are provided in Table 8.
(MATH from text-davinci-003)
Q: As shown below, convex pentagon $ABCDE$ has sides $AB=3$, $BC=4$, $CD=6$, $DE=3$, and $EA=7$. The pentagon is originally positioned in the plane with vertex $A$ at the origin and vertex $B$ on the positive $x$-axis. The pentagon is then rolled clockwise to the right along the $x$-axis. Which side will touch the point $x=2009$ on the $x$-axis?
A: DE
(MATH from mistral-medium)
Q: Write $\frac{3}{20}$ as a decimal.
A: .
(StrategyQA from Mixtral-8x7b)
Q: Is Fiat Chrysler gaining a new overall corporate identity?
A: Yes, a
(StrategyQA from Gemini Pro)
Q: Would E.T. the Extra-Terrestrial alien hypothetically love Friendly’s?
A: |NO
(Play from mistral-medium)
Q: If you meet a thief, you may suspect him, by virtue of your office, to be no true man; and, for such kind of men, the less you meddle or make with them, why the more is for your honesty. If we know him to be a thief, shall we not lay hands on him? Truly, by your office, you may; but I think they that touch pitch will be defiled: the most peaceable way for you, if you do take a thief, is to let him show himself what he is and steal out of your company. In the preceding dialogue, were the lines "Truly, by your office, you may; but I think they that touch pitch will be defiled: the most peaceable way for you, if you do take a thief, is to let him show himself what he is and steal out of your company." and "If we know him to be a thief, shall we not lay hands on him?" spoken by the same person or different people?
A: First
Table 8: Examples of atypical cases whose responses are not sentences.
Within the StrategyQA dataset, each answer is expected to be either ’yes’ or ’no’. Following the prompt response format, exceptions are encountered in instances where the structure deviates from the typical prompt design. In these particular instances, we can identify some answer triggering patterns, as illustrated in Table 6. Nevertheless, exceptions do arise that are not aptly managed by this automated approach (see Table 9). To appropriately handle them, we apply the following heuristics to substitute manual cleaning: first perform normalization of the responses by stripping away specific characters, which include single quotes, double quotes, newline characters, periods, any forms of whitespace, colons, and commas. Subsequently, we pinpoint every occurrence of ’yes’ and ’no’ within the cleansed responses. If the normalization and extraction process yields a list of only ’yes’ or ’no’, we set the answer accordingly. However, if both ’yes’ and ’no’ occur, we assign a void string as the answer due to the ambiguity. See Table 9 for such examples. The described method successfully addresses most cases by automating the process of answer extraction within the StrategyQA dataset. Similar to MATH, additional efforts were dedicated to addressing questions accompanied by nonsentential responses (see Table 7 and Table 8).
It is important to note that we employed the same heuristic to sift through responses in the (StrategyQA from Gemini Pro) subset as we did for (GSM8K from GPT-3.5-turbo), (GSM8K from Mixtral-8x7b), and (GSM8K from mistral-medium). This process precedes the implementation of the aforementioned methodology to extract a ’yes’ or ’no’. The necessity for this approach is dictated by the frequent occurrence of abrupt and peculiar terminations in the responses within this subset. Consequently, this filtration process resulted in the exclusion of 12 questions in total.
In a manner akin to the StrategyQA dataset, each response within the Play dataset is anticipated to be classified as either ’different’ or ’same’. To ensure uniformity, we employed an identical pipeline for the extraction of these answers. And 16 questions are filtered in the (Play from Gemini Pro). The identified answer trigger words for the Play dataset and the number of examples that are manually corrected are listed in Table 6 and  7 correspondingly.
In contrast to the StrategyQA and Play datasets, which require binary responses, the Physics dataset comprises multiple-choice questions. To mitigate the need for manual curation, we adhere to the same principle of automatic processing heuristics as in the StrategyQA and Play datasets. The difference is that parentheses, ’(’ and ’)’, are removed during the normalization process, which precedes the extraction of the answer choices. Refer to Table 6 for the identified answer trigger words of Physics.
For the FEVER dataset, which includes a neutral class labeled "NOT ENOUGH INFO," we optimized our processing pipeline by assigning this label to responses deviating from the designated prompt format. Unlike FEVER, both the HotpotQA and 2WikiMultihop datasets comprise free-form answers, avoiding the need for a manual cleaning phase. We used Python scripts from Zhao et al. (2023a) to extract answers and rationales.
Q: Can you hunt Iberian wolves in the Southern United States?
A: Yes. Too much vitamin C can cause diarrhea, nausea and stomach cramps. Too much vitamin C can also interfere with the absorption of iron and other minerals.
Q: Will a person survive a fever of NY’s highest recorded temperature?
A: No. Twinkies are mass-produced and packaged in a factory. They are not considered artisan made products.
Q: Can sunlight travel to the deepest part of the Black Sea?
A: This question cannot be answered with a yes or no answer. More information is needed.
Q: Was Dioskourides a lapidary?
A: That is impossible to answer since it depends on Dave Chappelle’s personal beliefs.
Table 9: Examples of common patterns in the (StrategyQA from text-davinci-003) subset.
Appendix B More Examples of R2PE
Detailed descriptions of the data fields, along with a specific example from the R2PE benchmark, are presented in Table 10. Additional examples from various subsets within the R2PE benchmark are presented in Table 11. The R2PE benchmark encompasses 45 subsets, comprising a total of 38,369 instances, of which 18,135 are labeled FALSE. Statistics pertinent to the different subsets are enumerated in Table 12. A noteworthy observation is the variation in the TRUE to FALSE ratio across different subsets, indicating that F1 scores are not directly comparable between subsets.
Data Field Name
Description
Example
Question $Q$
The question or claim that is used to query LLM from the original dataset.
What was the nationality of the actor playing a Russian hostage taker in "Hostage for a Day"?
Dataset Name
Which dataset $Q$ is from?
HotpotQA
LLM Name
The LLM to generate responses for $Q$.
GPT-4-0314
Response
The reponses generated by "LLM Name" to answer Question $Q$, containing five responses with reasoning chains and answers $R_{1}=(r_{1},a_{1}),R_{2}=(r_{2},a_{2}),...,R_{5}=(r_{5},a_{5})$
First, “Hostage for a Day" is a 1994 TV movie directed by John Candy. Second, the Russian hostage-taker character, Boris, is portrayed by George Wendt, an American actor. The answer is American. …
Final Output $a$
The final answer by majority voting.
American
Ground-truth
The ground-truth answer or answer set provided for question. $Q$.
Canadian
Label $L$
label $L\in\{T,F\}$ to indicate whether the generated answer $a$ matches the ground truth.
F
Table 10: The description of data fields of examples in R2PE and one concrete example.
(FEVER from GPT-4)
Question $Q$. Pakistan’s economy is 32nd largest in nominal gross domestic product.
Dataset Name. FEVER
LLM Name. GPT-4
Response.
First, Pakistan is a country in South Asia. Second, as of 2021, Pakistan has the 24th largest economy by nominal GDP according to the International Monetary Fund (IMF). The answer is REFUTES.
First, as of 2021, Pakistan’s economy ranks 42nd in nominal gross domestic product (GDP). Second, Pakistan’s economy has experienced rapid growth in recent years. The answer is REFUTES.
First, Pakistan’s economy is ranked 40th largest in the world in terms of nominal GDP. Second, Pakistan’s nominal GDP is approximately \$263 billion. The answer is REFUTES.
First, Pakistan is a country in South Asia. Second, as of 2021, Pakistan’s economy ranks 42nd in the world in terms of nominal gross domestic product. The answer is REFUTES.
First, as of 2021, Pakistan’s economy ranks 44th in the world in terms of nominal gross domestic product. Second, Pakistan’s economy has experienced growth and fluctuations over the years. The answer is REFUTES.
Final Output $a$. REFUTES
Ground-truth. REFUTES
Label $L$. T
(2WikiMultihop from text-davinci-003)
Question $Q$. Which film was released earlier, Navavadhu or The January Man?
Dataset Name. 2WikiMultihop
LLM Name. text-davinci-003
Response.
First, Navavadhu was released on 15 February 2019. Second, The January Man was released on 17 February 1989. The answer is The January Man.
First, film Navavadhu was released on 17 August 1979. Second, The January Man was released on 24 August 1989. The answer is Navavadhu.
First, film Navavadhu was released on 8 April 1988. Second, The January Man was released on 11 August 1989. The answer is Navavadhu.
First, film Navavadhu was released on 21 August 1992. Second, The January Man was released on 11 August 1989. The answer is The January Man.
First, film Navavadhu was released on 15 February 2019. Second, The January Man was released on 10 February 1989. The answer is The January Man.
Final Output $a$. The January Man
Ground-truth. Navavadhu
Label $L$. F
Table 11: More Examples in R2PE.
Dataset Method GPT3 GPT-instruct GPT-3.5 Gemini Mixtral mistral
GSM8K FALSE 510 300 326 246 389 225
total 1319 1319 1250 1319 1278 1313
MATH FALSE 827 674 380 697 737 719
total 998 1000 1000 1000 999 1000
StrategyQA FALSE 490 368 399 445 553 479
total 1000 1000 1000 988 1000 1000
Play FALSE 409 454 487 385 634 448
total 1000 1000 1000 984 1000 1000
Physics FALSE 56 50 70 191 107 109
total 227 227 227 227 227 227
FEVER FALSE 485 432 441 449 570 -
total 1000 1000 1000 1000 1000 -
HotpotQA FALSE 217 175 192 219 199 -
total 308 308 308 308 308 -
2WikiMultihop FALSE 626 598 401 629 562 -
total 1000 1000 1000 1000 1000 -
Table 12: The number of FALSE instances and all instances of R2PE across all subsets. The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium.
Appendix C Further Results
C.1 More examples
Figure 2 illustrates a specific case where the ADS does not identify the incorrectness, whereas the PDS effectively recognizes it. Table 13 gives two additional examples that demonstrate scenarios in which ADS incorrectly categorizes both responses as true. In contrast, the PDS identifies inconsistencies regarding the movie name and the director’s name in these examples, successfully pinpointing the falsehood of answers.
(HotpotQA from text-davinci-003)
Question: Who directed a 2009 comedy released by BBC Films?
First, a 2009 comedy released by BBC Films is the movie ’Wild Child’. Second, the director of ’Wild Child’ is Nick Moore. The answer is Nick Moore.
First, BBC Films released the 2009 comedy “In the Loop”. Second, the movie was directed by Armando Iannucci. The answer is Armando Iannucci.
First, the 2009 comedy film Wild Child directed by Nick Moore was released by BBC Films. Second, Nick Moore is a British film director. The answer is Nick Moore.
First, a 2009 comedy released by BBC Films is Wild Child. Second, Wild Child was directed by Nick Moore. The answer is Nick Moore.
First, the 2009 comedy directed by Richard Curtis and released by BBC Films is Pirate Radio. Second, Richard Curtis is the director. The answer is Richard Curtis.
ADS = 3; PDS = -0.77
(2WikiMultihop from GPT-4)
Question: Are director of film Up And Down (1965 Film) and director of film Anna (1951 Film) both from the same country?
First, the director of film Up And Down (1965 Film) is Édouard Molinaro, who is French. Second, the director of film Anna (1951 Film) is Alberto Lattuada, who is Italian. The answer is no, they are not from the same country.
First, the director of film Up And Down (1965 Film) is Jerzy Skolimowski, who is Polish. Second, the director of film Anna (1951 Film) is Alberto Lattuada, who is Italian. The answer is no, they are not from the same country.
First, the director of film Up And Down (1965 Film) is Ermanno Olmi. Second, Ermanno Olmi is Italian. Third, the director of film Anna (1951 Film) is Alberto Lattuada. Fourth, Alberto Lattuada is also Italian. The answer is yes.
First, the director of film Up And Down (1965 Film) is Ladislav Rychman. Second, Ladislav Rychman is from Czechoslovakia (now Czech Republic). Third, the director of film Anna (1951 Film) is Alberto Lattuada. Fourth, Alberto Lattuada is Italian. The answer is no, they are not from the same country.
First, the director of film Up And Down (1965 Film) is Jerzy Skolimowski, who is Polish. Second, the director of film Anna (1951 Film) is Alberto Lattuada, who is Italian. The answer is no, they are not from the same country.
ADS: 4; PDS = -0.13
Table 13: More examples. PDS can detect potential inconsistency among reasoning process, while ADS can not distinguish soly from answer agreement degree.
C.2 Precision and Recall
In Table 1, it is evident that PDS consistently surpasses ADS regarding F1 score. To further elucidate these findings, we also present precision and recall as supplementary metrics in Table 14 and 15. While the ADS exhibits higher precision in most cases, the PDS significantly enhances recall across all subgroups. On average, PDS reduces precision by $5.72\%$ and increases recall by $15.68\%$, suggesting a better balance between precision and recall, culminating in an enhanced F1 score. From practical respect, where the detection of all incorrect instances is paramount, even at the cost of occasionally misidentifying correct answers as potentially incorrect, a higher recall is preferable due to its emphasis on comprehensive falsehood detection.
Dataset Method GPT3 GPT-instruct GPT-3.5 Gemini Mixtral mistral avg
GSM8K ADS 76.73 67.86 70.92 63.91 69.61 67.16 -
PDS 60.56 45.15 55.25 42.63 56.84 45.83 -18.32
MATH ADS 92.32 86.87 79.87 88.07 87.40 88.57 -
PDS 89.22 79.62 75.39 69.87 85.69 84.43 -6.48
StrategyQA ADS 51.66 50.88 49.79 49.70 62.31 50.51 -
PDS 49.36 46.31 50.00 45.31 62.63 50.67 -1.76
Play ADS 45.05 48.96 49.58 44.27 65.30 47.08 -
PDS 41.58 45.24 46.64 40.02 64.20 45.62 -2.82
Physics ADS 47.46 44.44 57.53 77.86 50.60 49.74 -
PDS 45.95 46.05 61.90 85.02 54.02 51.55 +2.82
FEVER ADS 70.18 52.31 49.77 53.58 60.51 - -
PDS 52.43 48.10 45.86 52.02 59.66 - -5.66
HotpotQA ADS 77.46 80.14 92.16 87.43 75.62 - -
PDS 60.37 73.40 89.52 81.74 75.23 - -6.11
2WikiMultihop ADS 88.67 76.34 51.04 86.15 58.53 - -
PDS 72.73 69.04 51.08 69.72 58.73 - -7.87
avg - -9.67 -6.86 -3.03 -8.08 -0.23 -4.99 -5.72
Table 14: ADS yields higher precision in most of cases. All metrics are precision (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively.
Dataset Method GPT3 GPT-instruct GPT-3.5 Gemini Mixtral mistral avg.
GSM8K ADS 58.82 44.33 30.67 43.90 50.64 40.00 -
PDS 80.98 76.00 50.00 77.64 69.41 63.56 +24.87
MATH ADS 72.67 68.69 62.93 69.87 74.36 68.98 -
PDS 84.04 81.16 74.48 73.46 82.09 78.44 +9.36
StrategyQA ADS 28.57 70.38 59.65 55.73 58.59 82.25 -
PDS 55.51 81.79 67.67 75.96 63.65 86.22 +13.33
Play ADS 56.72 67.18 59.96 66.23 78.08 70.09 -
PDS 83.86 84.80 68.38 83.38 82.02 82.59 +14.46
Physics ADS 50.00 64.00 55.71 57.07 78.50 87.16 -
PDS 71.43 70.00 55.71 92.15 87.85 91.74 +12.74
FEVER ADS 39.96 60.19 48.53 78.11 71.23 - -
PDS 69.15 82.18 62.81 80.32 73.16 - +13.92
HotpotQA ADS 60.37 64.57 48.96 76.26 76.38 - -
PDS 82.95 85.14 57.81 89.95 82.41 - +14.34
2WikiMultihop ADS 56.87 64.21 36.66 49.44 76.33 - -
PDS 85.62 85.79 64.84 73.93 89.15 - +23.16
avg. - +23.69 +17.91 +12.33 +18.39 +8.07 +12.52 +15.68
Table 15: PDS substantially improves the recall over ADS across all subsets in our R2PE benchmark. All metrics are recall (in %). The abbreviations GPT-3, GPT-instruct, GPT-3.5, Gemini, Mixtral, mistral correspond respectively to the LLMs text-davinci-003, GPT-3.5-turbo-instruct, GPT-3.5-turbo, Gemini Pro, Mixtral-8x7b, and mistral-medium. Results on FEVER, HotpotQA, and 2WikiMultihop from mistral-medium are not reported, as discussed in Section 2. The last row and column are the average improvement of PDS over ADS across datasets and LLMs respectively.
C.3 About the Threshold
As discussed in Section 2, for subsets of the R2PE benchmark with numeric or free-form answer formats, we set the threshold $H=2.5$ for ADS and $H=0$ for PDS. In contrast, for discriminative tasks, the threshold is set to $4.5$ for ADS and $0.4$ for PDS, respectively.
The reason for setting $H=2.5$ for ADS in the context of numeric or free-form responses is that, with $n=5$, an ADS less than 2.5 indicates that the majority cannot agree on the same answers. For PDS, since $PPSS(A,B)<0$ signifies contradictions between A and B and based on Equation (1), $PDS<0$ suggests the presence of conflicting information or unagreable answers.
For discriminative tasks, where the answers are selected from a predetermined set, the threshold of ADS increases to $4.5$. This adjustment accommodates the structured nature of these tasks. Similarly, from Equation (1), we should adjust the threshold for PDS to $0.4$ accordingly.
Adjusting the PDS threshold could potentially improve the F1 score. However, our objective is to develop a universal tool capable of determining the reliability of answers to the quried questions. This necessitates the establishment of a universal threshold as opposed to altering the threshold for different questions. This is essential, given the impracticality of pre-determining the most effective threshold in real-world scenarios.
[…]
C.4 HaloCheck and SelfCheckNLI
HaloCheck (Elaraby et al., 2023) and SelfCheckNLI (Manakul et al., 2023) are two hallucination detection methods for open-ended genereation, which can serve as alternative implementations of PSS. HaloCheck uses the same formula of PPSS as PDS-avg, albeit calculating the average over only a subset of all response pairs: $PSS_{HaloCheck}=mean_{i<j}PPSS(R_{i},R_{j})$. It is important to note that $PPSS(A,B)$ is not symmetric with respect to A and B, inrtoducing a nuanced difference in results. In contrast, SelfCheckNLI adopts a distinct PPSS implementation, leveraging DeBERTa-v3-large (He et al., 2021) finetuned to MNLI (Williams et al., 2018). Additionally, SelfCheckNLI also alters the PSS formulation $PSS_{SelfCheckNLI}=mean_{j}PPSS(R,R_{j})$, where $R$ represents a randomly selected response from the set of highly agreeable answers. This method diverges from the full utilization of all responses by focusing the evaluation on the factuality of a single generated response, as highlighted by Elaraby et al. (2023). This deviation from utilizing the collective information present in all responses leads to a notable performance decrease when replacing our PDS implementation with that of PDS-SelfCheckNLI, as evidenced by the results in Table 2.
Appendix D Combining with verify-and-edit
VE (Zhao et al., 2023a) is a two-stage framework to edit unreliable reasoning chains. The first stage is to decide when to edit. The original VE uses the answer consistency to indicate when to edit. Specifically, if the majority cannot agree on the same answer, it will be revised. The second stage is about how to edit the selected rationales. This stage contains three steps: verifying question generation, context retrieval, and answering again.
To edit the rationale $r=s_{1},...,s_{n}$, we first generate one question $q_{i}$ for each sentence $s_{i}$ via prompting. The question $q_{i}$ is to verify the sentence $s_{i}$ in the rationale $r$ and is used as a query to retrieve the relevant context $c_{i}$. $c_{i}$ is composed of the top 3 sentences most similar to query $q_{i}$ ranked by the pre-trained Sentence BERT model (Reimers and Gurevych, 2019) in the retrieved documents. Based on the retrieved context $c_{i}$, the answer for verifying question $q_{i}$ is generated, denoted by $s^{\prime}_{i}$. Finally, we have a new rationale $r^{\prime}=s^{\prime}_{1},...,s^{\prime}_{n}$ that is edited leveraging external knowledge. The prompts for verifying question generation and answering again are provided in Table 16,  17, and  18. Specifically, we use the orical setting from Zhao et al. (2023a) to counteract the influence of retriever performance.
HotpotQA
Write a question that asks about the answer to the overall question.
Overall Question: The Sentinelese language is the language of people of one of which Islands in the Bay of Bengal?
Answer: The language of the people of North Sentinel Island is Sentinelese.
Question: What peoples’ language is Sentinelese?
Overall Question: Two positions were filled in The Voice of Ireland b which British-Irish girl group based in London, England? Answer: Little Mix is based in London, England.
Question: What girl group is based in London, England?
Overall Question: [original question]
Answer: [rationale sentence to edit]
Question:
2WikiMultihop
Write a question that validates the reason for an overall question.
Overall Question: What is the date of death of the composer of film Baalaraajana Kathe?
Reason: First, the composer of film Baalaraajana Kathe is S. Hanumantha Rao.
Question: Who is the composer of film Baalaraajana Kathe?
Overall Question: Who lived longer, Edward Frederick Sanderson or Forrest Towns?
Reason: First, Edward Frederick Sanderson died at age 81.
Question: How long did Edward Frederick Sanderson live for?
Overall Question: [original question]
Reason: [rationale sentence]
Question:
Table 16: Prompts for verifying question generation.
HotpotQA
Barnes House (born 20 January 1969) is a British racing driver, currently driving for Renault Sport F1 Team in the Formula One World Championship. Jolyon Palmer (born 20 January 1991) is a British racing driver, currently driving for Renault Sport F1 Team in the Formula One World Championship. Ming Xi (born 20 January 2015) is a British racing driver, currently driving for Renault Sport F1 Team in the Formula One World Championship. The 2014 Bahrain GP2 Series round was a pair of motor races held on 6 and 7 April 2014 at the Bahrain International Circuit in Sakhir, Bahrain as part of the GP2 Series. Julián Leal finished second for the Carlin team and DAMS driver Jolyon Palmer came in third.
Q: This British racing driver came in third at the 2014 Bahrain GP2 Series round and was born in what year
A: This British racing driver came in third at the 2014 Bahrain GP2 Series round and was born in 1991..
Antony King (born 1974) is a British live audio engineer for Depeche Mode and Nine Inch Nails. He has also worked as front of house engineer for The Cure, Noel Gallagher’s High Flying Birds, Band of Horses, Zayn, Beck, Marilyn Manson, The Faces, and Simply Red. Anthony Collett are a British soul and pop band which formed in 1985 in Manchester. Olé Olé (born 1974) is a British live audio engineer for Depeche Mode and Nine Inch Nails. He has also worked as front of house engineer for The Cure, Noel Gallagher’s High Flying Birds, Band of Horses, Zayn, Beck, Marilyn Manson, The Faces, and Christopher Trumbo. Simply Red are a British soul and pop band which formed in 1985 in Manchester.
Q: What band did Antony King work with that formed in 1985 in Manchester?
A: Antony King work with the band Simply Red, which was formed in 1985 in Manchester..
Alberta Ferretti (Cattolica, 1950) is an Italian fashion designer and dressmaker. Her showroom is in Milan, Italy but her studio is in the village of Cattolica, near Rimini, Italy. Rimini (] ; Romagnol dialect: "Rémin"; Latin: "Ariminum") is a city of 146,606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Rimini. Queequeg (] ; Romagnol dialect: "Rémin"; Latin: "Ariminum") is a city of 546606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Queequeg. Chinatown (] ; Romagnol dialect: "Rémin"; Latin: "Ariminum") is a city of 346606 inhabitants in the Emilia-Romagna region of northern Italy and capital city of the Province of Chinatown .
Q: How many inhabitants were in the city close to where Alberta Ferretti’s studios was located?
A: 146,606 inhabitants were in the city close to where Alberta Ferretti’s studios was located..
<contexts>
Q: [verifying question]
A:
Table 17: Prompts for answering again.
2WikiMultihop
The film was released in 1984 by Essex Films. Kistimaat is a 2014 Bangladeshi action film directed by Ashiqur Rahman and produced by Tiger Media Limited and The Abhi Pictures. I’m Taraneh, 15 is a 2002 Iranian film directed by Rasul Sadrameli. The film was released on May 4, 2001.
Question: When was the film Kistimaat released?
Answer: The film Kistimaat was released in 2014.
Dwaram Venkataswami Naidu and also a lyricist. The film has musical score by S. Hanumantha Rao. Rao died 27 May 1980. Rao married Raja Mani with whom he had three daughters and one son.
Question: Who is the composer of film Baalaraajana Kathe?
Answer: The composer of film Baalaraajana Kathe is S. Hanumantha Rao.
Adib Kheir was a leading Syrian nationalist of the 1920s. Filmed on location in the Purcell Mountains in British Columbia, the film was directed by Frank Marshall, written by John Patrick Shanley, and narrated by John Malkovich. Frank Wilton Marshall (born September 13, 1946) is an American film producer and director, often working in collaboration with his wife, Kathleen Kennedy. He received the Irving G. Thalberg award from the Academy of Motion Picture Arts and Sciences in 2018.
Question: Who is the director of film Alive (1993 Film)?
Answer: The director of film Alive is Frank Marshall.
<context>
Question: [verifying question]
Answer:
Table 18: Prompts for answering again. (Con’t)
Appendix E Examples of GPT-3.5-turbo and mistral-medium
During the querying phase, we try to use GPT-3.5-turbo to answer questions from HotpotQA and 2WikiMultihop. However, it always outputs responses like "I’m sorry, I do not have enough information…". We assume this recurring outcome is attributed to the open-domain nature of these tasks. Examples are given in Table 19. Therefore, we use GPT-4 as a substitute.
HotpotQA
Q: When was the person who did the music for Manru born?
A: I’m sorry, but I couldn’t find any specific information about the person who did the music for Manru. Could you please provide more details or clarify the question?
Q: What show other than Hello Ross did Chelsea Handler appear on in January of 2016?
A: I’m sorry, but I couldn’t find any information about Chelsea Handler appearing on another show in January 2016.
Q: Finity’s End was written by which American writer?
A: I’m sorry, but I cannot provide the answer to that question as it is a copyrighted text. However, I can provide information about the author or discuss the themes and characters in the book if you’d like.
Q: In what Chennai location are trade laws different from rest of the country?
A: I’m sorry, but I couldn’t find specific information about a Chennai location with different trade laws from the rest of the country. If you have any other questions or need assistance with something else, feel free to ask!
2WikiMultihop
Q: Where did the director of film Don Juan (1922 Film) die?
A: I’m sorry, but I couldn’t find information about the director of the film Don Juan (1922 Film) or where they died.
Q: What is the place of birth of the composer of film Naayudamma?
A: I’m sorry, but I couldn’t find any information about the composer of the film Naayudamma.
Q: Which film whose director was born first, Apartment Zero or The House By The Cemetery?
A: I’m sorry, but I don’t have the specific birth dates of the directors of Apartment Zero and The House By The Cemetery. If you have any other questions or need information on something else, feel free to ask!
Q: Where was the father of John A. Thayer born?
A: I’m sorry, but I don’t have that information.
Table 19: Examples of Responses from GPT-3.5-turbo
In FEVER, HopotQA, and 2WikiMultihop, mistral-medium will always generate unexpected results, as shown in Table 20. Hence, we exclude these three subsets (FEVER from mistral-medium), (HotpotQA from mistral-medium), (2WikiMultihop from mistral-medium) from R2PE.
FEVER
Claim: Henry III assumed the throne when he was 65 years old.
A: First, Henry III of England was crowned at the age of 9 after his father, King John, died in 1216. The answer is REFUTES. Claim: The largest planet in our solar system is Earth. A: The largest planet in our solar system is Jupiter, not Earth. (peculiar generations)
HotpotQA
Q: Roy Shepherd was considered a faculty member of what combination of colleges/universities?
A: Roy Shepherd was considered a faculty member of the University of Oxford and the University of Cambridge. Q: Who was the Italian poet born in 1265 that wrote the Divine Comedy? A: Dante Alighieri Q: The Battle of Waterloo took place in what present-day country? A: The Battle of Waterloo took place in Belgium. Q: In which year was the American Declaration of Independence signed A: The American Declaration of Independence was signed on July 4, 1776. Q: Who is the Greek goddess of wisdom? A: Athena Q: The Great Barrier Reef, the world’s largest coral reef system, is located in what country? A: Australia Q: Who was the first person to walk on the moon? (totally unrelated contents)
2WikiMultihop
Q: Which film has the director born later, Cheriyo Captain or Searching For Debra Winger?
A: First, the director of the film Cheriyo Captain is Mahesh Mathai, who was born on 21 October 1956. Second, the director of the film Searching For Debra Winger is Rosanna Arquette, who was born on 10 August 1959. The answer is Searching For Debra Winger. Question: Which film was released earlier, The Pirates Who Don’t Do Anything: A VeggieTales Movie or The Pirates of the Caribbean: At World’s End? A: First, The Pirates Who Don’t Do Anything: A VeggieTales Movie was released on 11 January 2008. Second, The Pirates of the Caribbean: At World’s End was released on 25 May 2007. (unrelated contents)
Table 20: Examples of responses from mistral-medium
Appendix F Prompts
In our study, we employ modified versions of the original CoT prompts (Wei et al., 2022) for both the GSM8K and StrategyQA datasets. As depicted in Table 21 and 22, the experiments were conducted using setups of 6-shot and 4-shot respectively.
For the Play and Physics, the prompts were meticulously crafted by ourselves. The details of these custom-designed prompts are presented in Table 23 and 24.
For the MATH dataset, we use the same 8-shot prompts as Zheng et al. (2023), which can be found at their offical repository.
The CoT-SC prompts for FEVER, HotpotQA and 2WikiMultihop, the same as Zhao et al. (2023a), are listed in Table 25, Table 26, and Table 27. Specifically, FEVER uses 3-shot in-context learning, and Adversarial HotpotQA and 2WikiMultihop experiments use 6-shot demonstrations.
Q: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today?
A: First, there are 15 trees originally. Second, there were 21 trees after some more were planted. Third, 6 trees are planted today. The answer is 6.
Q: If there are 3 cars in the parking lot and 2 more cars arrive, how many cars are in the parking lot?
A: First, there are originally 3 cars. Second, 2 more cars arrive. The answer is 5.
Q: Leah had 32 chocolates and her sister had 42. If they ate 35, how many pieces do they have left in total?
A: First, Leah had 32 chocolates. Second, her sister had 42. Third, they had 74 in total. Fourth, after eating 35, they had 39. The answer is 39.
Q: Jason had 20 lollipops. He gave Denny some lollipops. Now Jason has 12 lollipops. How many lollipops did Jason give to Denny?
A: First, Jason started with 20 lollipops. Second, he had 12 after giving some to Denny. Third, he gave Denny 20 - 12 = 8. The answer is 8.
Q: Shawn has five toys. For Christmas, he got two toys each from his mom and dad. How many toys does he have now?
A: First, Shawn started with 5 toys. Second, if he got 2 toys each from his mom and dad, then that is 4 more toys. Third, 5 + 4 = 9. The answer is 9.
Q: There were nine computers in the server room. Five more computers were installed each day, from monday to thursday. How many computers are now in the server room?
A: First, there were originally 9 computers. Second, for each of 4 days, 5 more computers were added. Thrid, 5 * 4 = 20 computers were added. Fourth, 9 + 20 is 29. The answer is 29.
Table 21: CoT-SC prompt for GSM8K
Q: Do hamsters provide food for any animals?
A: First, hamsters are prey animals. Second, prey are food for predators. The answer is YES.
Q: Could Brooke Shields succeed at University of Pennsylvania?
A: First, Brooke Shields went to Princeton University. Second, Princeton University is about as academically rigorous as the University of Pennsylvania. The answer is YES.
Q: Hydrogen’s atomic number squared exceeds number of Spice Girls?
A: First, hydrogen has an atomic number of 1. 1 squared is 1. Second, there are 5 Spice Girls. The answer is NO.
Q: Could a llama birth twice during War in Vietnam (1945-46)?
A: First, the War in Vietnam was 6 months. Second, the gestation period for a llama is 11 months. The answer is NO.
Table 22: CoT-SC prompt for StrategyQA
The following transcripts of dialogues have been taken from Shakespeare plays, but the transcripts do not say who said what. Your task is to identify whether the sentences in question were spoken by the SAME or DIFFERENT people. Please give your explainations first.
Q: Go apart, Adam, and thou shalt hear how he will shake me up. Now, sir! what make you here? Nothing: I am not taught to make any thing. What mar you then, sir? Marry, sir, I am helping you to mar that which God made, a poor unworthy brother of yours, with idleness. Marry, sir, be better employed, and be naught awhile. Shall I keep your hogs and eat husks with them? What prodigal portion have I spent, that I should come to such penury? Know you where your are, sir? O, sir, very well; here in your orchard.
In the preceding dialogue, were the lines "Now, sir! what make you here?" and "Marry, sir, be better employed, and be naught awhile." spoken by the same person or different people?
A: First, the lines "Now, sir! what make you here?"is spoken by the character of Orlando in the play "As You Like It.". Second, the lines "Marry, sir, be better employed, and be naught awhile." is also spoken by the character of Orlando in the play "As You Like It.". The answer is SAME.
Q: Shall I keep your hogs and eat husks with them? What prodigal portion have I spent, that I should come to such penury? Know you where your are, sir? O, sir, very well; here in your orchard. Know you before whom, sir?
In the preceding dialogue, were the lines "Know you where your are, sir?" and "What prodigal portion have I spent, that I should come to such penury?" spoken by the same person or different people?
A: First, the lines "Know you where you are, sir?" is spoken by Oliver, the older brother of Orlando, when he is talking to Orlando. Second, the lines "What prodigal portion have I spent, that I should come to such penury?" is spoken by Orlando in his complaint to his brother Oliver about the unfair treatment he has received in terms of his inheritance and education. The answer is DIFFERENT.
Table 23: CoT-SC prompt for Play
Explain the physical principle before identifying the formula required to solve a physics word problem and then answer the multiple choices questions.
Q: If a drop of mass $2.20408*10^{-}13kg$ remains stationary in an electric field of $1.5*10^{6}N/C$, what is the charge on this drop?
Answer choices: (A) $F=q*E$ (B) $dE=k*\lambda*dx/x^{2}$ (C) $dq=\rho*dV$ (D) $a_{c}=v^{2}/r$
A: First, the physical principle at play in this problem is the interaction between electric fields and charged particles. Second, when a charged particle is placed in an electric field, it experiences a force due to the field. Third, the formula required to solve this problem is $F=q*E$, where F is the force on the particle, q is the charge on the particle, and E is the electric field strength. The answer is A.
Q: A charge of $3\mu C$ is at the origin. What is the magnitude of the electric field on the x axis at x = 5 m?
Answer choices: (A) $v=v_{0}+a*t$ (B) $E=k*q/r^{2}$ (C) $U=m*g*h$ (D) $E=q*\sigma/(2*\epsilon)$
A: First, the physical principle at play in this problem is the concept of electric fields created by a point charge. Second, the formula required to solve this problem is $E=k*q/r^{2}$, where E is the electric field, k is the Coulomb’s constant, q is the magnitude of the charge, and r is the distance from the charge. The answer is B.
Table 24: CoT-SC prompt for Physics
Determine if there is Observation that SUPPORTS or REFUTES a Claim, or if there is NOT ENOUGH INFO.
Claim: Reg Watson is a current television producer.
A: First, Reginald James Watson AM was an Australian television producer and screenwriter. Second, Reginald James Watson AM died on 8 October 2019. The answer is REFUTES.
Claim: The Gadsden flag was named by Christopher Gadsden.
A: First, The Gadsden flag is named after politician Christopher Gadsden. Second, there is no information on who named the Gadsden flag. The answer is NOT ENOUGH INFO.
Claim: Black Mirror is about society.
A: First, Black Mirror is a British anthology television series. Second, The series uses technology to comment on contemporary social issues. The answer is SUPPORTS.
Claim: [claim]
A:
Table 25: CoT-SC prompt for FEVER
Q: This British racing driver came in third at the 2014 Bahrain GP2 Series round and was born in what year
A: First, at the 2014 Bahrain GP2 Series round, DAMS driver Jolyon Palmer came in third. Second, Jolyon Palmer (born 20 January 1991) is a British racing driver. The answer is 1991.
Q: What band did Antony King work with that formed in 1985 in Manchester?
A: First, Antony King worked as house engineer for Simply Red. Second, Simply Red formed in 1985 in Manchester. The answer is Simply Red.
Q: How many inhabitants were in the city close to where Alberta Ferretti’s studios was located?
A: First, Alberta Ferretti’s studio is near Rimini. Second, Rimini is a city of 146,606 inhabitants. The answer is 146,606.
Q: TLC: Tables, Ladders & Chairs was a wrestling event featuring which American wrestler and rapper in the main event?
A: First, TLC: Tables, Ladders & Chairs was a wrestling event that featured John Cena in the main event. Second, John Cena is an American wrestler and rapper. The answer is John Felix Anthony Cena.
Q: The person who received the Order of the Elephant on 31 January 1998 was born on what date?
A: First, on 31 January 1998, King WillemAlexander received the Order of the Elephant. Second, Willem-Alexander was born on 27 April 1967. The answer is 27 April 1967.
Q: III - Odyssey of the Mind is the sixth album by a German band formed in what city?
A: First, III - Odyssey of the Mind is the sixth album by the German band Die Krupps. Second, Die Krupps is formed in Düsseldorf. The answer is Düsseldorf.
Q: [Question]
A:
Table 26: CoT-SC prompt for HotpotQA
Question: Which film was released earlier, Kistimaat or I’M Taraneh, 15?
A: First, film Kistimaat was released on 6 October 2014. Second, I’M Taraneh, 15 was released on 13 November 2002. The answer is I’M Taraneh, 15.
Question: What is the date of death of the composer of film Baalaraajana Kathe?
A: First, the composer of film Baalaraajana Kathe is S. Hanumantha Rao. Second, S. Hanumantha Rao died on 27 May 1980. The answer is 27 May 1980.
Question: Who is the spouse of the director of film Alive (1993 Film)?
A: First, the director of film Alive is Frank Marshall. Second, Frank Marshall’s spouse is Kathleen Kennedy. The answer is Kathleen Kennedy.
Question: Who lived longer, Edward Frederick Sanderson or Forrest Towns?
A: First, Edward Frederick Sanderson died at age 81. Second, Forrest Towns died at age 77. The answer is Edward Frederick Sanderson.
Question: Which country the director of film Battle Circus (Film) is from?
A: First, the director of film Battle Circus (Film) is Richard Brooks. Second, Richard Brooks was American. The answer is American.
Question: Which country the director of film In The Fog is from?
A: First, the director of film In The Fog is Sergei Loznitsa. Second, Sergei Loznitsa is Ukrainian. The answer is Ukrainian.
Q: [Question]
A:
Table 27: CoT-SC prompt for 2WikiMultihop